OpenStudy (anonymous):

In right triangle ABC, C is the right angle and sin A = sqrt3/2. what is the value of cscB?

OpenStudy (anonymous):

hmm....... Cosines..... so here we go! can i get a figure

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] take this prime formula!

OpenStudy (anonymous):

\[a^2\over c^2 \]+\[ b^2\over c^2 \] But a2/c2 = (sin A)2, and b2/c2 = (cos A)2. In order to reduce the number of parentheses that have to be written, it is a convention that the notation sin2 A is an abbreviation for (sin A)2, and similarly for powers of the other trig functions. Thus, we have proven that \[\sin^2 A + \cos^2 A = 1\]

OpenStudy (anonymous):

This problem can be ground out and solved, but I think the real issue here is being able to recognize the special values of sin x that come from a 30-60-90 triangle. sin A = sqrt3/2 indicates a sixty degree angle at A; in turn, this indicates a 30 degree angle at C, which means sin B=1/2, and csc B =2. This can all be solved, but as a trig student, you should know these special values, and be able to make the connection when you encounter them.